OpenStudy (anonymous):

16. What is the equation of the line that passes through the points (–2, 3) and (2, 7)? (4 points) A. x – y = –1 B. x – y = –2 C. x – y = –5 D. x – y = –6

OpenStudy (eric_d):

find the slope first

OpenStudy (mathstudent55):

The equation of the line through points \((x_1, y_1) \) and \((x_2, y_2) \) is: \(y - y_1 = \dfrac{y_2 - y_1}{x_2 - x_1} (x - x_1) \). You are given two points. Let \(x_1 = -2\), \(y_1 = 3\), \(x_2 = 2\), and \(y_2 = 7\). Then substitute those values in the equation above and find the equation of the line.

OpenStudy (eric_d):

then... y=x+b

OpenStudy (mathstudent55):

@eric_d You mean y = mx + b, right?

OpenStudy (eric_d):

?